OpenStudy (anonymous):

How much interest will be paid in 1 year for a loan of $3600 at 8 1/2% simple interest? will give medal

OpenStudy (jdoe0001):

\(\bf i=prt \\ \quad \\ p=3,600\qquad t=1 \\ \quad \\ r=8\frac{1}{2}\%\implies \cfrac{8\frac{1}{2}}{100}\implies \cfrac{\frac{17}{2}}{100}\implies \cfrac{17}{2}\cdot \cfrac{1}{100}\implies 0.085\)

OpenStudy (texaschic101):

Interest = principle * rate * time I = PRT P = 3600 T = 1 R = 8 1/2% = 8.5% = .085 now sub I = 3600 (.085)(1) can you finish ?

OpenStudy (texaschic101):

jdoe...you type fast...lol